Which is an equation of the line that passes through the point (2, 0) and is parallel to the graph of y = –4x + 1?
Answer by FrankM(1040)

The full equation you give is of the form y=mx+b where m is the slope.
Here, it's -4
Simply put, parallel lines have the same slope.
So we need to create a line with slope -4 that passes through (2,0) so we use the form:
y-0=-4(x-2)
ANS y=-4x+8